### v2

# Edit As You Wish: Video Caption Editing with Multi-grained User Control

Abstract.
Automatically narrating videos in natural language complying with user requests, i.e. Controllable Video Captioning task, can help people manage massive videos with desired intentions. However, existing works suffer from two shortcomings: 1) the control signal is single-grained which can not satisfy diverse user intentions; 2) the video description is generated in a single round which can not be further edited to meet dynamic needs. In this paper, we propose a novel Video Caption Editing (VCE) task to automatically revise an existing video description guided by multi-grained user requests. Inspired by human writing-revision habits, we design the user command as a pivotal triplet {operation, position, attribute} to cover diverse user needs from coarse-grained to fine-grained. To facilitate the VCE task, we automatically construct an open-domain benchmark dataset named VATEX-EDIT and manually collect an e-commerce dataset called EMMAD-EDIT. We further propose a specialized small-scale model (i.e., OPA) compared with two generalist Large Multi-modal Models to perform an exhaustive analysis of the novel task. For evaluation, we adopt comprehensive metrics considering caption fluency, command-caption consistency, and video-caption alignment. Experiments reveal the task challenges of fine-grained multi-modal semantics understanding and processing. Our datasets, codes, and evaluation tools are ready to be open-sourced.
1. Introduction
The proliferation of videos on the Internet heralds the era of video-dominated media. Video captioning, i.e. automatically describing videos using natural language, has been a prevalent task to assist people in comprehending and managing massive videos. However, conventional video captioning systems (Venugopalan et al., 2015; Yu et al., 2017) tend to generate intention-agnostic descriptions, ignoring the various demands of different users. Therefore, a new task branch, namely controllable video captioning (Chen et al., 2019; Deshpande et al., 2019; Yuan et al., 2020; Liu et al., 2021), has been proposed to integrate user intention as a control signal to guide the description generation.
Although controllable video captioning has great potential in practical applications, existing works have two non-negligible drawbacks. First, they all employ fixed control signals that can only express single-grained controls, such as Part-of-Speech(POS) (Wang et al., 2019a) for structure control, or specified object tags (Liu et al., 2021) for semantic control. These single-grained controls can not satisfy flexible and diverse user demands. Second, they are single-round controls that generate a video description once which can not be further revised. Whereas iteratively revising sentences until the ideal texts is a natural process for humans (Du et al., 2022). Imagine a real-world scenario, such as E-commerce product promotion, where sellers upload product videos with descriptions to attract customers. There is a good chance that automated video descriptions fail to highlight the seller’s preferences. As a result, sellers need to further improve the descriptions by themselves, which is time-consuming and labor-intensive, especially when facing massive long videos.
We propose a novel Video Caption Editing (VCE) task to automate the video description editing process. The task aims to edit an existing video description conditioned on user commands and video content. As depicted in Figure 1, the inputs of VCE task consist of a video, a reference description, and a user control. It outputs an edited video description based on the user command as a control signal. The reference caption can be initialized using the last edited output sentence which can thus enable multi-round modification. The VCE task can facilitate personalized video description generation by fulfilling miscellaneous demands from different users or dynamic demands from the same user.
In the VCE task, how to define the user command to cover multi-grained requests is crucial. Inspired by the human writing-revision habits, we unify the user edit commands into a triplet format {operation, position, attribute} (depicted in Figure 2) for three advantages. Firstly, it condenses the core elements in an editing operation. Secondly, it can accommodate two prevalent front-end interface signals including natural language and editing trajectories from tablet computers. Finally, the different combinations of three elements in the triplet can cover multi-grained user commands from coarse-grained control (e.g. sentence length change) to fine-grained control (e.g. insert new details in the specific position), as illustrated in Table 1.
We collect two novel benchmark datasets named VATEX-EDIT and EMMAD-EDIT to support the exploration of the VCE task. The VATEX-EDIT dataset is automatically constructed from a large-scale video-text dataset VATEX (Wang et al., 2019b) in the open domain. Meanwhile, to close the gap between research advances and real-life applications, we manually collect an e-commerce editing dataset called EMMAD-EDIT, which is more challenging from two aspects: 1) longer videos (average 27.1 seconds) and longer captions (average 100 words); 2) external domain knowledge needed to generate product-oriented video descriptions. Based on the two benchmark datasets, we propose a specialized model, namely OPA, that converts the command triplet into a textual token sequence to alleviate heterogeneity among multi-grained commands. We demonstrate the feasibility of utilizing a unified framework to handle seven types of user commands. Moreover, we adopt two generalist Large Multimodal Models (LMMs) as a comparison to gain an in-depth understanding of the characteristics and challenges of the VCE task.
The main contributions of this paper are four-fold. 1) To the best of our knowledge, we are the first to propose the VCE task to achieve multi-round editing and design the user command as a triplet format to express multi-grained user requests. 2) We build two benchmark datasets from different domains, including the general domain (VATEX-EDIT) and commercial domain (EMMAD-EDIT), to facilitate the investigation of the VCE task. 3) We develop an evaluation suite to assess the edited video description based on caption fluency, command-caption consistency, and video-caption alignment. 4) We propose a unified specialist framework OPA and adapt two generalist LMM methods to initially tackle the task, followed by a comprehensive analysis.
2. Related Work
Controllable Video Captioning. Video captioning (Ramanishka et al., 2016; Venugopalan et al., 2015; Tang et al., 2021; Yu et al., 2017; Aafaq et al., 2019; Luo et al., 2020; Pan et al., 2020; Shetty and Laaksonen, 2016; Shi et al., 2019) is a challenging cross-modal task to automatically describe the visual contents of a video in natural languages. In order to satisfy the varied pragmatic interests of different users, controllable video captioning (Chen et al., 2019; Deshpande et al., 2019; Yuan et al., 2020; Liu et al., 2021) has been a newly prevalent task. It aims to derive video descriptions conditioned on a predefined control signal, e.g. visual object tags. Wang et al. (2019a) introduce Part-of-Speech(POS) information as guidance and Yuan et al. (2020) directly utilize an exemplar sentence. Their goal is to generate descriptions with desired syntactic structures. Meanwhile, other works aim to control sentence semantics. Chen et al. (2019) proposes a topic-guided model to generate topic-oriented descriptions. Liu et al. (2021) focus on producing object-oriented sentences controlled by multiple user-interested objects. However, the above endeavors all generate a sentence once and can’t be edited dynamically. Besides, their designed control signals are single-grained which can not cover flexible user intentions. Instead, we define a novel VCE task that can revise a description in multiple rounds covering multi-grained user demands.
Command Notation Demand Unified Input Control opera. pos attr ✓ add, - , - expand description [ADD] A group of girls is playing a game. ✓ ✓ add, pos, - expand description at specified positions [ADD] A group of girls is [MASK] playing a game. ✓ ✓ add, - , attr add specified attributes in description [ADD] field, hockey; A group of girls is playing a game. ✓ ✓ ✓ add, pos, attr add specified attributes at specified positions [ADD] field, hockey; A group of girls is [MASK] playing a game. ✓ del, - , - shorten description [DEL] A group of girls is on the field playing a game. ✓ ✓ del, pos, - shorten description at specified positions [DEL] A group of girls is on (the filed) [MASK] playing a game. ✓ ✓ del, - , attr delete specified attributes from description [DEL] field, group; A group of girls is on the field playing a game.
Image Caption Editing. Conventional image captioning (Vinyals et al., 2015; Xu et al., 2015; Liu et al., 2016; Stefanini et al., 2022; Li et al., 2016; Liu et al., 2018; Yao et al., 2023; Chen et al., 2017; Yang et al., 2023a) generates the description for images from scratch which may lead to factual mistakes. Sammani and Elsayed (2019) firstly define the image caption editing task that modifies an existing caption (a.k.a. reference caption) conditioned on the image content to obtain more accurate descriptions. Sammani and Melas-Kyriazi (2020) further propose a novel EditNet framework to achieve interactive and adaptive edits. Yuan et al. (2023) design an adaptive text-denoising network to alleviate the semantic gap between input images and reference sentences. The above works all edit image captions implicitly. Wang et al. (2022) propose the explicit image caption editing task to make the modification process more explainable and efficient. In summary, these caption editing works can only correct the wrong content in the reference captions and ignore specific edit intentions of different users. In this paper, we integrate multi-grained user commands into the video description editing process.
Large Multimodal Models. Recent months have witnessed the tremendous success of Large Language Models (LLMs) (Chowdhery et al., 2022; Ouyang et al., 2022; Zeng et al., 2022; Touvron et al., 2023b, a; OpenAI, 2022, 2023a) towards artificial general intelligence. Further, Large Multimodal Models (LMMs) (Dai et al., 2023; Liu et al., 2023; Bai et al., 2023; Ye et al., 2023a; Cha et al., 2023; Li et al., 2023a; Zhang et al., 2023; Luo et al., 2023; Lin et al., 2023; Ren et al., 2023) endow LLMs with the visual understanding ability by incorporating vision backbones (Sun et al., 2023; Radford et al., 2021; Li et al., 2022b). Existing LMMs primarily bifurcate into two categories: Image Large Language Models (ImgLLMs) (Dai et al., 2023; Liu et al., 2023; Bai et al., 2023; Ye et al., 2023a; Cha et al., 2023) and Video Large Language Models (VidLLMs) (Li et al., 2023a; Zhang et al., 2023; Lin et al., 2023; Luo et al., 2023; Ren et al., 2023). The standard architecture of an LMM comprises a vision backbone for encoding images or videos, a projector (Li et al., 2023b; Cha et al., 2023) to condense and translate visual embeddings into textual semantic space, and an LLM to process all multimodal contexts. As the VCE task involves capturing video semantics, understanding textual user controls, and enabling multi-grained text editing, it can serve as a new touchstone for LMMs.
3. Video Caption Editing Task
3.1. Task Definition
Given a video and a reference caption , the VCE task aims to generate an edited caption according to the user edit command . The edited caption should satisfy the constraints of , and . Given a ground truth caption , the maximum likelihood estimation (MLE) training objective of VCE task can be formulated as:
The reference caption can be initialized with the output caption from the last round of editing. It is also possible to start the editing process using a auto-generated sentence or human-written one as the reference. Due to the reference caption input setting, the VCE task can naturally achieve an interactive editing process with successive editing rounds, which is in line with human writing habits (Du et al., 2022). The interactive and multi-round revisions can help produce descriptions with higher user satisfaction.
Dataset Vision #Videos/Images #Editing instances VTime Edit Dist Vocab Train Val Test Train Val Test COCO-EE (Wang et al., 2022) Image 52,587 3,055 2,948 97,567 5,628 5,366 - 10.3 9.7 10.9 11,802 Flickr30K-EE (Wang et al., 2022) Image 29,783 1,000 1,000 108,238 4,898 4,910 - 7.3 6.2 8.8 19,124 VATEX-EDIT Video 25,467 2,935 5,867 784,805 91,513 181,638 10.0 14.4 16.0 11.9 21,634 EMMAD-EDIT Video 16,176 5,418 5,502 47,569 15,914 16,169 27.1 91.3 93.7 17.8 44,725
3.2. User Edit Command
It is not trivial to define flexible edit commands in the VCE task to meet various realistic user needs. We observe that natural language and writing-revision traces are two natural interactive modes. The former can be received from keyboards or speech converters, while the latter conveniently expresses user intentions with the prevalence of tablets and wireless stylus pens111https://www.apple.com/apple-pencil/. A command representation compatible with the above two signals is important and meaningful. In this paper, we propose a novel command representation in a triplet format {operation, position, attribute}, where operations control the overall description editing, positions specify the editing locations, which could affect the syntax of sentences, and attributes guide the editing operation to control the semantic contents of descriptions.
We define the atomic edit operations as add and delete, considering that the replace editing can be decomposed into the two atomic operations (i.e. first delete then add). Meanwhile, position and attribute in the triplet are optional, therefore, as shown in Table 1, seven specific commands222Note that we omit the command “del, pos, attr, delete attributes at specified positions”, as it can be covered by “del, pos, -, delete description at specified positions”. via different combinations of operation, position, and attribute elements in the triplet can cover multi-grained realistic demands from coarse-grained (global) controls to fine-grained (local) controls. The designed triplet command can be flexibly obtained by processing the inputs from front-end interfaces including natural language and writing-revision traces (details in Appendix E). In the following method and experiments sections, we perform video description editing directly based on the triplet command.
4. Data Collection
To faciliate the novel VCE task, we automatically construct an open-domain dataset VATEX-EDIT, and manually annotate an E-commerce dataset EMMAD-EDIT. Table 2 displays the overall data statistics. Compared to prior image caption editing datasets such as COCO-EE and Flickr30K-EE, our new datasets present several distinct advantages: 1) more challenging with the video input and lengthier captions; 2) more diverse encompassing open-domain and e-commerce data; and 3) larger in scale. Specific annotated data instances are illustrated in Figure 3.
4.1. VATEX-EDIT Construction
It is challenging to construct data samples for the VCE task from scratch, which needs a quadruple (video, command, reference caption, edited caption) data, abbreviated as (). To mitigate the difficulty, we build the VATEX-EDIT dataset by expanding the widely-used video captioning dataset VATEX (Wang et al., 2019b), which has high-quality caption annotations for each video.
We sample an annotated caption of a video as the reference caption, and the next goal is to construct the command and the related edited caption. In general, we aim to construct related (command, edited caption) samples including: 1) coarse-grained length-control commands referring to the global add or delete edits that result in length changes, and 2) finer-grained attribute-related commands to achieve add or delete attributes on the reference caption.
Coarse-grained length-control commands. For add operation, we directly select a longer caption with significant length differences from the rest annotations as the edited caption. The delete operation does the exact opposite. Considering the delete operation can be easily achieved without video content referring, we replace the reference caption with negative captions that have partially misaligned semantics with the video content. Such updated quadruples require models to prioritize removing visually irrelevant content from the reference, which makes the delete more challenging.
Fine-grained attribute-related commands. We construct attribute samples in a “degradation” manner, that is firstly detecting attribute words in the reference caption and then removing them to obtain the edited caption . We utilize Spacy Syntactic Dependency333https://spacy.io/ and Semantic Role Labeling (Shi and Lin, 2019) to achieve noun, verb or modifier attributes detection and removal in while maintaining the fluency of to get . After degradation, we can obtain quadruples for the del, -, attr command. We exchange the reference caption and the edited caption to obtain the opposite add command. The position information can be naturally recorded to support position-related commands.
Statistics and analysis. As shown in Table 2, compared with existing image editing datasets (Wang et al., 2022), VATEX-EDIT has two salient features: large-scale and diverse. Figure 4 visualizes the percentages of modifiers, nouns and verbs in the attribute. Our automatic construction strategy selects verbs as the dominant attributes because verbs are usually related to temporal visual semantics, which is also one of the core challenges of the video description task. The word cloud of specific attribute words shows the attribute diversity.
4.2. EMMAD-EDIT Collection
To satisfy realistic user-demand scenarios, we manually collect a high-quality dataset called EMMAD-EDIT in the E-commerce domain based on a Chinese E-commerce video captioning dataset E-MMAD (Zhang et al., 2022). The E-MMAD dataset consists of product videos with advertising video descriptions, and additional structure information. Given a product-oriented video and an original video description , we recruit crowd workers to annotate three types of edited descriptions as follows.
Simplify original captions to the target length while maintaining sentence fluency and coherence according to the video content. To ensure the challenge of the VCE task, we require that the length of original sentences should be reduced by at least 20%.
Delete specific attributes. It aims to select multiple significant attribute words/phrases from and remove the attribute-related content to get a new caption . The attributes can be nouns, verbs, or modifiers. To ensure the semantic coherence of , workers are allowed to modify other parts of following the “minimal editing” principle.
Delete abstract attributes. We further consider deleting abstract attributes that do not directly appear in . For example, deleting “Time and Seasons” needs to locate season-related content such as “spring” and “summer”. It is more challenging to edit with abstract attributes and also more down-to-earth since user intentions may be vague.
Statistics and analysis. To ensure annotation quality, extra workers further check the annotated cases. Table 2 shows the specific data statistics. EMMAD-EDIT has two remarkable characteristics, i.e. long videos and long descriptions. The average video length is 27.1 seconds and the average description length (specified in Figure 5) is around 100 words. We believe the challenging EMMAD-EDIT dataset will promote new technologies for the VCE task.
5. Methodology
In this section, we begin by introducing how to transform the triplet control into a unified textual sequence. Subsequently, we explore three approaches for the VCE task to facilitate a comprehensive comparison. We propose the Operation-Position-Attribute (OPA) model as a small-scale specialist. Additionally, we utilize an Image Large Language Model (ImgLLM) pipeline, and an end-to-end Video Large Language Model (VidLLM) to observe the performance of large multimodal models. Lastly, we develop an evaluation protocol for the novel task.
5.1. Input Format Design
We first integrate the seven specific edit commands introduced in Table 1 into a unified format to achieve multi-grained control.
The main challenge is the heterogeneity among the three elements of command, including operation, position, and attribute. On the one hand, operations and attributes change the textual semantics while positions mainly influence sentence syntax. On the other hand, attributes are specific textual words while positions are absolute position indexes.
To tackle the above challenges, we unify the input format as a textual token sequence. As shown in Table 1, we define two special tokens, [ADD] and [DEL], to represent different add or delete operations. Attribute words are naturally text tokens. For position, we put special tokens [MASK] in the reference caption to indicate the absolute position indexes. For example, a positioned reference caption “A group of girls is [MASK] playing a game” guides the model to generate new details between words “is” and “playing”. Finally, we concatenate the operation token, the attribute words, and a positioned reference caption as a control sequence to guide the model for description generation. Table 1 visualizes the input control sequences under seven specific commands when the reference caption is “A group of girls is playing a game”.
5.2. OPA: A Small-Scale Model as the Specialist
We construct a small-scale encoder-decoder Transformer architecture, i.e. multi-modal BART (Lewis et al., 2020), to achieve the video description editing task under the guidance of processed control sequences. We utilize the pre-trained BART weights and endow it with the multi-modal ability to understand video content. The specific architecture is depicted in Appendix A.
Input Representation. Given a video , a reference caption , and a triplet command , we first extract frame-level visual features and map them to the same dimension as textual embedding. We denote the input attributes as and the indicated position index as . Taking the most fine-grained command “add specified attributes at specified positions” as an example, the concatenated control sequence for the command is defined as . Using special tokens to separate each part, it is formulated as:
where the positioned reference caption is formulated as:
Finally, we input the visual features and the textual control sequence embedding to the Transformer encoder. If the position is empty in the command, we input the original reference caption . When the attribute is empty in the command, we set as an empty set.
Leverage Pre-trained Knowledge. The overall training objective as formulated in Section 3.1 is to generate an edited description conditioned on the video features and the control sequence. It is worth noting that we keep the [MASK] token consistent with the same token in the Text Infilling pre-trained task of BART to leverage the intrinsic pre-training textual ability.
5.3. LMMs as Contrastive Generalists
Large multimodal models integrate the advantages of visual understanding and remarkable natural language processing abilities (e.g., text editing) from LLMs. It is significant to probe their performance on the VCE task. Consequently, we explore two typical branches of LMMs including an ImgLLM pipeline and an end-to-end VidLLM.
ImgLLM Pipeline. We utilize the InstructBLIP (Dai et al., 2023) as the ImgLLM. Nevertheless, ImgLLM can only handle images or a single video frame as visual input. To adapt the ImgLLM to the VCE task, we combine InstructBLIP with ChatGPT (OpenAI, 2022). In this way, InstructBLIP transforms visual semantics at the frame level into textual context, while ChatGPT consolidates all textual task context and achieves caption editing. Specifically, we extract frames from a given video and utilize InstructBLIP to produce detailed visual descriptions for each frame. The frame descriptions with the VCE task definition, task guidelines, and in-context demonstrations (Dong et al., 2022) of the relative command type will be combined as the final prompt to the ChatGPT. Instruction details are provided in Appendix A.
End-to-end VidLLM. As VidLLM can handle video tasks directly, we employ the VideoChat (Li et al., 2023a) as an end-to-end LMM solution. Specifically, we reformat the VATEX-EDIT and EMMAD-EDIT datasets into question-answer chat samples (refer to Appendix A for specifics) and conduct further instruct-tuning on VideoChat-7B using two datasets respectively.
5.4. Evaluation Suite
How to evaluate the novel VCE task is another noteworthy challenge. Conventional video captioning tasks adopt widely-used captioning metrics such as BLEU4, METEOR, and CIDEr. However, these reference-based metrics only measure the consistency between generated captions and ground-truth annotations, which are insufficient. In this paper, we evaluate the VCE task from three aspects: 1) fluency, 2) controllability, and 3) text-video alignment.
Fluency. Following the previous work (Wang et al., 2022), we adopt widely-used BLEU4 (Papineni et al., 2002) and ROUGE-L (Lin, 2004) metrics to measure the overall generation quality. We also use the Perplexity (PPL) (Jelinek et al., 1977) metric that reflects the grammatical correctness and semantic meaningfulness.
Controllability. Measuring whether an edited caption strictly follows control signals is important for the VCE task. Inspired by previous work (Du et al., 2022) , we first utilize the SARI (Xu et al., 2016) metric to measure the overall edit quality, i.e. the consistency between expected-to-edit and actually-edited spans. Moreover, we design three breakdown metrics namely Length Accuracy, Attribute Accuracy and Position Accuracy to measure whether the edited caption satisfies the {operation, position, attribute} triplet control. Concretely, Len-Acc reflects the length change accuracy. Attr-Acc checks the appearance of commanded attribute words. Pos-Acc evaluates whether the model inserts/removes content in the specified positions.
Text-Video alignment. The VCE task inherently requires the alignment between edited descriptions and the given video. We use EMScore (Shi et al., 2022) to calculate the semantic similarity between edited captions and videos. It focuses on both coarse-grained similarity (video-sentence) and fine-grained similarity (frame-word).
6. Experiments
| Model | Controllability | Fluency | Vision Align | ||||||
| Len-Acc | Attr-Acc | Pos-Acc | SARI | BLEU4 | ROUGE-L | PPL | EMScore | ||
| 1 | Multimodal BART | - | - | - | 49.7 | 48.0 | 62.0 | 73.9 | 28.7 |
| 2 | 97 | - | - | 52.1 | 49.6 | 63.2 | 70.6 | 28.7 | |
| 3 | 97 | 93 | - | 53.8 | 52.3 | 65.7 | 72.7 | 28.7 | |
| 4 | OPA | 98 | 93 | 99 | 53.9 | 52.2 | 65.6 | 72.8 | 28.7 |
| 5 | 98 | 82 | 97 | 52.6 | 50.5 | 64.4 | 77.2 | 28.7 | |
| 6 | 3 Single-grained Models | 96 | 69 | 99 | 53.3 | 51.5 | 64.6 | 72.8 | 28.6 |
Command Controllability Fluency Vision Align Len-Acc Attr-Acc Pos-Acc SARI BLEU4 ROUGE-L PPL EMScore 1 add, - , - 57 - - 26.2 62.1 73.9 23.7 44.7 2 add, pos, - 75 - 59 27.0 83.6 90.3 25.1 44.9 3 add, - , attr 80 74 - 31.7 84.7 89.9 26.6 45.2 4 add, pos, attr 92 70 85 32.9 88.1 93.3 25.5 44.8 5 del, - , - 100 - - 33.5 66.8 73.8 30.5 44.6 6 del, pos, - 99 - - 30.7 83.9 90.6 28.8 44.6 7 del, - , attr 100 93 - 33.6 75.2 83.6 28.9 44.7 8 Overall 84 79 72 30.5 74.4 81.8 26.6 44.8
6.1. Implementation Details
We implement the small-scale OPA model based on Huggingface Transformers library (Wolf et al., 2020). The default setting is initialized by the . We get video frames using fps=1. For the VATEX-EDIT dataset in English, we adopt BLIP (Li et al., 2022a) ViT-B/16 to extract frame features. The max frame sequence N is set to 20. For the EMMAD-EDIT dataset in Chinese, we initialize our model with the Chinese BART. We adopt CN-CLIP (Yang et al., 2022) ViT-B-16 to extract video frame-level features. The max frame sequence N is set to 30 and the max decoding length is set to 150. For training, we use AdamW (Loshchilov and Hutter, 2017) with a learning rate of 1e-5 and optimize for 20 epochs with a batch size of 20. During inference, we set beam size as 5. The ImgLLM pipeline utilizes the identical frame number N as the OPA model. This pipeline doesn’t involve any training. We choose one in-context learning sample for every command type integrated into the ChatGPT prompt. For VideoChat model, we set frame number N as 8 to fit its default setting. We conduct further instruct-tuning on the official 7B checkpoints with batch size 64.
6.2. Compare Specialist and Generalist Models
Figure 6 shows the overall performance of three baselines on the VATEX-EDIT and EMMAD-EDIT datasets. Interestingly, overall performances are divergent across the two datasets. On the large-scale open-domain VATEX-EDIT dataset, the small-scale specialist OPA model with only 150M parameters outperforms the LMM approaches. It suggests that with sufficient training instances (784,805 samples in VATEX-EDIT), a small specialized model has the potential to perform more effectively and efficiently.
On the e-commerce EMMAD-EDIT dataset, the LMM methods achieve higher scores across most metrics. EMMAD-EDIT is more challenging because it requires domain knowledge, such as unseen product attributes and advertising description style, and has limited training data (refer to Table 2). Results show that ImgLLM with InstructBLIP and ChatGPT achieve highest Attr-Acc (87%) even without training. We argue that on this domain, generalist methods are more promising to leverage their intrinsic knowledge to edit product-related descriptions.
Despite performance advantages, the training cost and inference speed must be taken into account due to the booming number of videos. As illustrated in Figure 6 (right), compared to the VidLLM and ImgLLM, the OPA model demonstrates significant benefits over both VidLLM and ImgLLM in terms of lower training costs and faster inference speeds. Designing a model that balances performance with speed and cost represents a crucial trade-off.
Although both specialist and generalist models offer unique advantages, there remains considerable scope for further enhancements to develop an effective editing system, especially on the EMMAD-EDIT dataset. The Controllability (Pos-Acc 72%, Attr-Acc 87% and Len-Acc 94%) on the EMMAD-EDIT dataset is insufficient. Moreover, the alignment between video and caption, as indicated by the EMScore, requires significant improvement. In conclusion, the key of the VCE task lies in the combination of fine-grained video and user control understanding and precise text editing capabilities.
6.3. Further Task Analysis
We conduct further ablation studies on the small-scale OPA model to delve into a detailed analysis of the characteristics of the VCE task.
Increase control signals. We analyze the editing performance under different control signals in Table 3. The proposed OPA framework achieves high controllability accuracy (Len-Acc 98%, Attr-Acc 93%, and Pos-Acc 99%) while maintaining sentence quality. The first block (lines 1-4) shows the controllability accuracy and caption quality when progressively inputting more control signals into the model. With the increasing aspects of control signals, there is no decline in sentence fluency and text-vision alignment. It indicates that our model can edit the reference caption with reasonable syntactic and semantic changes under multi-aspect guidance. Line 5 shows the result of Pure Transformer trained from scratch. Without BART pre-trained parameters, the overall controllability and fluency metrics decrease (SARI from 53.9 to 52.6, BLEU4 from 52.2 to 50.5), which verifies the benefits of textual pre-training knowledge.
Unified framework vs separate models. To satisfy different-granularity commands, we compare the performances of training a unified OPA model vs. training multiple separate models in Table 3. In the 3 Single-grained Models setting (line 6), we train three models respectively to deal with three control granularities, i.e. {operation}, {operation, attribute}, and {operation, position, attribute}. The OPA model reaches a remarkably higher score on the Attr-Acc (93% vs 69%) with better SARI, BLEU4, and ROUGE-L. It demonstrates that our unified input design can alleviate the confusion and heterogeneity of multi-grained commands.
Difficulty level of various commands. Table 4 (lines 1-7) displays the performances of different commands, indicating their respective difficulty levels. Generally, the add operation proves to be more challenging than delete, primarily because it requires the constraint of video content. For the add operations, the finer the command granularity (lines 1-4), the higher the controllable and fluency scores. It reveals that when provided with more detailed control signals, the model can generate desired captions more easily.
Command Video EMScore SARI BLEU4 Overall ✗ 28.1 53.5 51.7 ✔ 28.7 (+0.6) 53.9 (+0.4) 52.2 (+0.5) del, - , - ✗ 27.0 39.4 12.1 ✔ 28.5 (+1.5) 40.4 (+1.0) 13.5 (+1.4) add, - , - ✗ 28.3 41.9 10.6 ✔ 29.0 (+0.7) 42.0 (+0.1) 11.1 (+0.5)
Effects of vision modality. We compare the model performance with and without video input in Table 5. Adding vision modality brings overall metric improvements since it provides visual semantics to guide edited video description generation. For del,-,- command, we especially construct challenging samples in which reference captions have misalignments with videos (Section 4.1). With the visual semantics, our model prioritizes removing the video-misalignment contents and achieving a higher EMScore (from 27.0 to 28.5). Similarly, add,-,- command requires enriching the original caption referring to the video content.
6.4. Quantitative Results
Multi-grained editing controls. Provided with various commands, the OPA model can output different edited descriptions to satisfy multi-grained user requests. As Figure 7 (a) shows, our OPA model successfully generates different desired descriptions given the same video, the same reference caption but different commands from coarse-grained (e.g. add, -, -) to fine-grained (e.g. add, , door).
Successive editing controls. The OPA model also supports interactive editing with successive controls in the VCE task, depicted in Figure 7 (b). The edited description can serve as the reference caption in the next round to make further editing to satisfy dynamic user demands.
| Dataset | Model | Control. | Fluency | Vision Align |
| EMMAD-EDIT | Trans. | 2.94 | 3.02 | 3.27 |
| OPA | 3.98 | 3.85 | 3.76 | |
| GT | 4.67 | 4.37 | 4.22 | |
| VATEX-EDIT | Trans. | 4.18 | 4.23 | 3.76 |
| OPA | 4.36 | 4.43 | 3.93 | |
| GT | 4.48 | 4.34 | 4.41 |
Human Evaluation. We further adopt human evaluation to assess the quality of edited video descriptions. We recruit 20 evaluators to score the generated descriptions. We randomly sample 200 test cases from VATEX-EDIT and 350 cases from EMMAD-EDIT respectively. During the evaluation, we randomly order the edited captions generated from Pure Transformer baseline, OPA, and groundtruths (GT). The evaluators are asked to rate each description from three aspects on a scale of 1 to 5 points. Table 6 shows the OPA model exceeds the controllable Transformer baseline in three aspects, especially the controllability.
7. Conclusion
We propose a novel multi-modal task named Video Caption Editing (VCE), which aims to automatically edit video descriptions under the guidance of multi-grained user commands. To satisfy diverse and varied user demands, we design the user control signal as a {operation, position, attribute} triplet to flexibly cover both coarse-grained and fine-grained controls. We collect two datasets named VATEX-EDIT and EMMAD-EDIT from different domains. We further employ comprehensive metrics to assess fluency, controllability, and vision-text alignment. Finally, we introduce a small specialized model called OPA, an ImgLLM pipeline, and an end-to-end VidLLM to dive into the task challenges and provide good starting points.
Limitations and Future Work. This paper primarily introduces appropriate baseline solutions for the VCE task, aiming to provide a thorough analysis. Nonetheless, it falls short of designing architectural innovations, leaving ample room for exploration in the future. Further insights into significant future directions are discussed in Appendix F.
Appendix A Model Details
A.1. Architecture of OPA Model
The overall architecture of our specialist model, i.e. OPA, is depicted in Figure 8.
A.2. Prompt of ImgLLM Pipeline
We build a ChatGPT pipeline with the visual expert model, i.e. InstructBLIP, to verify the vision-enhanced LLM performance on the VCE task. Specifically, we extract key frames from a given video and utilize InstructBLIP to produce detailed descriptions for each frame. We uniformly select 20 frames for VATEX-EDIT dataset and 30 frames for EMMAD-EDIT dataset, which is exactly the same frame number that our OPA model uses. These frame descriptions with timestamps help the ChatGPT to understand the exhaustive video content.
Designed Prompts. As Figure 10 illustrates, we conduct a well-designed prompt for ChatGPT including the task definition, helpful guidelines, an in-context learning (Dong et al., 2022) example, the video information, and the new case to be solved. With the input prompt, ChatGPT can output the desired edited caption with the required format.
In-Context Learning. Considering the differences between seven multi-grained commands, we manually select seven high-quality input-output demonstrations involving all command types. The command type of given examples will be aligned with the new case to fulfill better results. For example, if we want ChatGPT to edit a video description under the add, pos, attr command, we will give a matched example of add, pos, attr command to help it better solve the task.
A.3. Instruction-tuning Data of VidLLM
We convert the original samples from the VATEX-EDIT and EMMAD-EDIT datasets into an instruction-tuning format to facilitate the training of end-to-end video large language models. Figure 9 illustrates the converted data sample.
Appendix B VATEX-EDIT Dataset Details
B.1. Automatic Dataset Construction
We construct quadruples (video, command, reference caption, edited caption) according to two types of commands including coarse-grained length-control commands and fine-grained attribute-related commands. The overall procedure is depicted in Figure 11. We complement more details about constructing attribute-related commands.
Attribute-related commands. Our goal is to construct related (command, edited caption) samples for each (video, reference caption) to support attribute related commands in a “degradation” manner. The main challenges lie in two aspects: 1) extracting meaningful noun, verb, or modifier attributes in the reference caption and 2) deleting the attribute-related semantic spans while maintaining the fluency of rest content to get an attribute-removed caption . In detail, we adopt four steps as follows:
First, we use the Spacy444https://spacy.io/ syntactic dependency parser to build a textual dependency tree that contains the Part-of-Speech information and relationships between tokens. We select reasonable branches in the parsed tree as attributes and further prune the branch to ensure the fluency of the rest caption.
Second, we use a Semantic Role Labeling model (Shi and Lin, 2019) to analyze the semantic roles of each span in a sentence. It can help to better judge whether a parsed attribute span in the first step can be deleted or not, especially noun attributes.
Third, we merge the attributes which only modify one or two tokens to improve the task challenge, that is, the model may be required to edit with multiple attributes in one round.
Finally, considering the intrinsic error of the parsing model, we adopt a post-processing stage to filter low-quality sentences considering sentence fluency, edited token length, and attribute diversity. The ground-truth sentence fluency is further verified in Section B.2.
When the attribute-related spans are removed in a sentence, the edited positions can be naturally recorded. Through the above steps, we can get high-quality samples for the del, pos, attr commands. Meanwhile, reversed samples can be obtained for the add, pos, attr command by exchanging the reference caption and the edited caption. For these two fine-grained commands, the sentence length, structure, and semantics are controlled at the same time.
For commands add, -, attr and del, -, attr that omit positions, they mainly control the sentence length and semantics, not structure. We get relevant samples by relaxing the structure constraint based on the above “degradation” manner. In detail, we replace the edited caption by retrieving desired sentences satisfying both length and semantics (attributes) constraints from original annotated descriptions for the same video.
| Constructed | Annotated | |
| ChatGPT | 87% | 91% |
| Human | 4.40 | 4.37 |
B.2. VATEX-EDIT Test set Quality
In the VATEX-EDIT dataset, it is worth noting that only the “del, pos, -, shorten description at specified positions” command uses the auto-constructed sentences as ground-truth captions in order to control the sentence structure. Meanwhile, the data samples of other commands utilize human-annotated sentences from the original VATEX dataset as ground-truth captions. To assess the quality of the auto-constructed sentences, we evaluate the test set of VATEX-EDIT through human and ChatGPT evaluations. For the ChatGPT555https://openai.com/blog/chatgpt evaluation, we design suitable prompts (depicted in Figure 12) to guide ChatGPT to judge whether a sentence is fluent or not. We randomly sample 100 ground-truth sentences on the test set and calculate the fluency rate obtained by ChatGPT. For human evaluation, we recruit 20 crowd workers to rate the fluency score (ranging from 1 to 5) of 200 randomly sampled ground-truth sentences. As Table 7 shows, the automatically constructed ground-truth sentences of the “del, pos, -” command are as fluent as the human-annotated ground-truth sentences of other commands (87% vs 91% and 4.40 vs 4.37), which indicates the high quality of the VATEX-EDIT test set.
| Command | #Videos | #Editing instances | VTime | Edit Dist | Uni. Attrs | Vocab | ||||||
| Train | Val | Test | Train | Val | Test | |||||||
| add, - , - | 7,751 | 2,599 | 2,633 | 7,752 | 2,599 | 2,633 | 26.6 | 72.2 | 100.9 | 29.8 | - | 29,241 |
| add, pos, - | 3,221 | 1,077 | 1,094 | 3,221 | 1,077 | 1,094 | 26.3 | 91.6 | 99.4 | 8.7 | - | 18,640 |
| add, - , attr | 3,221 | 1,078 | 1,094 | 3,221 | 1,078 | 1,094 | 27.1 | 93.1 | 100.6 | 8.7 | 2,388 | 18,501 |
| add, pos, attr | 3,221 | 1,077 | 1,093 | 3,221 | 1,077 | 1,093 | 26.6 | 92.0 | 99.9 | 8.7 | 2,907 | 18,584 |
| del, - , - | 7,751 | 2,599 | 2,633 | 7,753 | 2,599 | 2,633 | 26.4 | 100.1 | 71.4 | 29.6 | - | 29,456 |
| del, pos, - | 3,221 | 1,078 | 1,095 | 3,221 | 1,078 | 1,095 | 27.1 | 102.1 | 86.1 | 17.2 | - | 18,734 |
| del, - , attr | 3,221 | 1,078 | 1,095 | 3,221 | 1,078 | 1,095 | 27.2 | 102.7 | 86.2 | 17.8 | 3,038 | 18,924 |
| 16,176 | 5,418 | 5,502 | 31,610 | 10,586 | 10,737 | 26.9 | 91.3 | 90.4 | 20.8 | 6,003 | 44,725 | |
| 15,955 | 5,328 | 5,432 | 15,959 | 5,328 | 5,432 | 26.8 | 90.9 | 100.9 | 11.4 | 648 | 44,347 | |
Appendix C EMMAD-EDIT Dataset Details
We manually collect the high-quality dataset EMMAD-EDIT in the E-commerce domain based on the Chinese E-commerce video captioning dataset E-MMAD. The data sample in E-MMAD dataset consists of a product video, an advertising video description, and additional information including video titles and structure attributes. It collects 120,984 videos with average duration of 30.4 seconds and the annotated Chinese description length is 67 words on average. The characteristics of long videos and long captions make it suitable for building a challenging VCE dataset. During annotation, we select data samples from E-MMAD dataset with relatively long descriptions. In addition to videos and descriptions, we also provide product structure information and video titles for reference.
C.1. EMMAD-EDIT dataset statistics
Table 8 shows breakdown data statistics of the EMMAD-EDIT dataset. It has overall 79,652 editing instances for 16,176 product videos. Diverse unique attributes and vocabulary also indicate data richness. We separate the abstract attribute-related data samples as an extra challenging subset. Considering realistic demands, we utilize all these data samples to construct “add, -, attr, add attributes in description” command cases. Note that in the Experiments section, we present the EMMAD-EDIT results training without the abstract subset data by default.
C.2. The Impact of Data Volume.
Considering the limited scale of manually collected data in EMMAD-EDIT, we analyze the results under different volume data with 4K, 8K, 12K samples. Figure 13 shows that a growing volume of data consistently increases the controllable scores. Breakdown analysis of multi-grained commands reveals that more challenging commands, e.g. add,-,-, require higher volume of training data samples to get desired performance.
C.3. Data Visualization
We manually collect data samples that support the editing of two types of attributes: specific and abstract. Specific attributes directly appear in the reference caption to support straightforward content editing such as “comfortable” and “fashion”. Meantime, abstract attributes are more high-level concepts that may consist of multiple specific attributes. For example, “Style”, “Target People”, and “Time and Seasons” are annotated abstract attributes in the dataset. The word clouds (shown in Figure 14) of specific and abstract attributes show the diversity of the EMMAD-EDIT dataset.
We visualize annotated samples of the EMMAD-EDIT dataset in Figure 15. Besides quadruples (video, command, reference caption, edited caption) that support video caption editing, we also provide additional product information such as structured information and the video title. The annotation interface during data construction is shown in Figure 16.
Appendix D Metric Details
To support evaluations in Chinese, we utilize a Chinese GPT-2 (Du, 2019) to calculate Chinese sentence likelihood and then obtain PPL scores. For EMScore, we replace the core vision-language alignment model EN-CLIP (Radford et al., 2021) with CN-CLIP (Yang et al., 2022).
D.1. Position Accuracy Design
We propose a novel Position Accuracy (Pos-Acc) metric to measure whether models insert/remove the content in the specified positions under fine-grained editing control. The challenge of calculating Pos-Acc is the misalignment between the reference caption and the edited caption . The two textual sentences have variable lengths and the edited operation may appear in the other positions to maintain the overall fluency. To tackle the above challenge, we propose a novel Dynamic Sequence Aligning (DSA) algorithm to align two variable-length textual sequences based on the absolute positions, inspired by classical Dynamic Time Warping (DTW) (Müller, 2007).
The POS-Acc metric can align two variable-length text sequences in positional indexes, resulting in related spans {} in aligned to [MASK] tokens in . We count the percentage of correct samples that insert/remove new content in given position .
As Figure 17 illustrates, we visualize two aligned cases in English and Chinese respectively obtained by the DSA algorithm.
Appendix E Conversion from Interface Signals to Triplet Format
Though the main focus of this paper lies in exploring the novel VCE task utilizing triplet commands, we also shed light on how to convert the prevalent interface signals, i.e. natural language and handwriting editing trajectories, to the triplet formatted controls.
E.1. From Natural Language to Triplets
To convert natural language signals, we can adopt fuzzy matching to recognize add or delete operations and use text parser tools to get specific semantic roles of a sentence. We can also leverage the ability of LLMs such as ChatGPT and LLaMA-2, which already have outstanding language understanding and summarization capabilities. Specifically, we can design suitable prompts to convert text sentences into triplets to meet the pre-defined requirements.
E.2. From Triplets to Natural Language
The triplet command in our annotated datasets can be easily converted to natural languages to support more diverse scenarios and applications. For example, using natural language to explore the capability of LLMs in the VCE task (A.2). Specifically, we design a series of templates shown in Table 9 to convert the triplet command to natural language. We will also release the two benchmark datasets with user commands both in the triplet and natural language format to benefit the community.
| Command | Conversion Template |
| del, - , attr | Delete contents about ’{}’ from the reference description. |
| add, pos, attr | Add contents about ’{}’ at [POS]. |
| add, - , attr | Add contents about ’{}’ to expand the reference description. |
| add, pos, - | Add video-related contents at [POS]. |
| del, pos, - | Contents have been deleted at [POS], please make the rest sentence fluent and coherent. |
| del, - , - | Shorten the reference description. |
| del, - , attr | Expand the reference description. |
E.3. From Handwriting-revision Traces to Triplets
The recent advancement of GPT-4Vision (GPT-4V) has shown its powerful capability of Visual Referring Prompting (Yang et al., 2023b). GPT-4V can well understand visual pointers (such as circles, arrows or traces) directly drawn on images, therefore, revealing a novel human-model interaction method called “visual referring prompting”. Combined with its accurate OCR capability, GPT-4V can serve as an ideal tool to input the handwriting editing traces as an image and convert it to the triplet control output, as illustrated in Figure 18.
Appendix F Future Directions
In this paper, we make the first attempt to propose the novel Video Caption Editing (VCE) task and collect two benchmark datasets to support it. We further propose a unified framework OPA for VCE and compare it with a ChatGPT pipeline. Based on the task, dataset and method foundation, we aim to make a good start for the community. There are various interesting aspects worth exploring in the future:
A versatile system for video captioning and editing. The dense annotation of the quadruple (video, command, reference caption, edited caption) in our dataset has the potential to support building a versatile system encompassing conventional video captioning, controllable video caption and video caption editing. For initialization, conventional video captioning can produce a generated description for a given video. subsequently, video caption editing can be utilized to update and revise the original description. When omitting the reference caption, the rest annotation (video, command, edited caption) can be adjusted to achieve controllable video captioning.
A more robust system for poor reference caption. In the VCE task, the reference caption can be the edited caption from the last round to fulfill multi-round editing. Furthermore, it can also be extended with human-written drafts or machine-generated captions. Under these circumstances, the robustness of the VCE system to low-quality reference captions should be taken into consideration.
The abstract-attribute subset of EMMAD-EDIT is under-explored. The abstract-attribute subset (details in Sec 4.2.) involves abstract attributes that do not directly appear in the reference caption. It requires models to understand and reason the video content at a higher semantic level. In the experiments, we only assess the performances of OPA and ChatGPT pipeline on the easier specific subset, whose performances are not yet so desirable, the abstract subset therefore remains a challenge for future exploration.
Serve as a touchstone for video large language models (VidLLMs). The recent emergence of VidLLMs such as GPT4-Vision (OpenAI, 2023b), VideoChat (Li et al., 2023a), and VideoChatGPT (Maaz et al., 2023) has ignited sparks for a generalist video assistant. However, existing research also points out their limitations for long video understanding (Jin et al., 2023), visual hallucination (Shukang et al., 2023) and multi-modal instruction following capability (Ye et al., 2023b). The VCE benchmark can serve as a new multi-modal evaluation for assessing both long video understanding and multi-grained text editing abilities.

### v3

@@
 Experiments reveal the task challenges of fine-grained multi-modal semantics understanding and processing.
-Our datasets, codes, and evaluation tools are ready to be open-sourced.
+Our datasets, codes, and evaluation tools are available at https://github.com/yaolinli/VCE.
 1. Introduction
@@
 2. Related Work
+Command Notation Demand Unified Input Control opera.
+pos attr ✓ add, - , - expand description [ADD] A group of girls is playing a game.
+✓ ✓ add, pos, - expand description at specified positions [ADD] A group of girls is [MASK] playing a game.
+✓ ✓ add, - , attr add specified attributes in description [ADD] field, hockey; A group of girls is playing a game.
+✓ ✓ ✓ add, pos, attr add specified attributes at specified positions [ADD] field, hockey; A group of girls is [MASK] playing a game.
+✓ del, - , - shorten description [DEL] A group of girls is on the field playing a game.
+✓ ✓ del, pos, - shorten description at specified positions [DEL] A group of girls is on (the filed) [MASK] playing a game.
+✓ ✓ del, - , attr delete specified attributes from description [DEL] field, group; A group of girls is on the field playing a game.
 Controllable Video Captioning.
@@
 Instead, we define a novel VCE task that can revise a description in multiple rounds covering multi-grained user demands.
-Command Notation Demand Unified Input Control opera.
-pos attr ✓ add, - , - expand description [ADD] A group of girls is playing a game.
-✓ ✓ add, pos, - expand description at specified positions [ADD] A group of girls is [MASK] playing a game.
-✓ ✓ add, - , attr add specified attributes in description [ADD] field, hockey; A group of girls is playing a game.
-✓ ✓ ✓ add, pos, attr add specified attributes at specified positions [ADD] field, hockey; A group of girls is [MASK] playing a game.
-✓ del, - , - shorten description [DEL] A group of girls is on the field playing a game.
-✓ ✓ del, pos, - shorten description at specified positions [DEL] A group of girls is on (the filed) [MASK] playing a game.
-✓ ✓ del, - , attr delete specified attributes from description [DEL] field, group; A group of girls is on the field playing a game.
 Image Caption Editing.
-Conventional image captioning (Vinyals et al., 2015; Xu et al., 2015; Liu et al., 2016; Stefanini et al., 2022; Li et al., 2016; Liu et al., 2018; Yao et al., 2023; Chen et al., 2017; Yang et al., 2023a) generates the description for images from scratch which may lead to factual mistakes.
+Conventional image captioning (Vinyals et al., 2015; Xu et al., 2015; Liu et al., 2016; Stefanini et al., 2022; Li et al., 2016; Liu et al., 2018; Yao et al., 2023; Yang et al., 2023a; Yao et al., 2022) generates the description for images from scratch which may lead to factual mistakes.
 Sammani and Elsayed (2019) firstly define the image caption editing task that modifies an existing caption (a.k.a.
@@
 (2022) propose the explicit image caption editing task to make the modification process more explainable and efficient.
-In summary, these caption editing works can only correct the wrong content in the reference captions and ignore specific edit intentions of different users.
+In summary, all these works can only correct the wrong content in the reference captions and ignore specific edit intentions of different users.
 In this paper, we integrate multi-grained user commands into the video description editing process.
@@
 Recent months have witnessed the tremendous success of Large Language Models (LLMs) (Chowdhery et al., 2022; Ouyang et al., 2022; Zeng et al., 2022; Touvron et al., 2023b, a; OpenAI, 2022, 2023a) towards artificial general intelligence.
-Further, Large Multimodal Models (LMMs) (Dai et al., 2023; Liu et al., 2023; Bai et al., 2023; Ye et al., 2023a; Cha et al., 2023; Li et al., 2023a; Zhang et al., 2023; Luo et al., 2023; Lin et al., 2023; Ren et al., 2023) endow LLMs with the visual understanding ability by incorporating vision backbones (Sun et al., 2023; Radford et al., 2021; Li et al., 2022b).
+Further, Large Multimodal Models (LMMs) (Dai et al., 2023; Liu et al., 2023; Bai et al., 2023; Ye et al., 2023a; Li et al., 2023a; Zhang et al., 2023; Lin et al., 2023; Ren et al., 2023) endow LLMs with the visual understanding ability by incorporating vision backbones (Sun et al., 2023; Radford et al., 2021; Li et al., 2022b).
 Existing LMMs primarily bifurcate into two categories: Image Large Language Models (ImgLLMs) (Dai et al., 2023; Liu et al., 2023; Bai et al., 2023; Ye et al., 2023a; Cha et al., 2023) and Video Large Language Models (VidLLMs) (Li et al., 2023a; Zhang et al., 2023; Lin et al., 2023; Luo et al., 2023; Ren et al., 2023).
-The standard architecture of an LMM comprises a vision backbone for encoding images or videos, a projector (Li et al., 2023b; Cha et al., 2023) to condense and translate visual embeddings into textual semantic space, and an LLM to process all multimodal contexts.
-As the VCE task involves capturing video semantics, understanding textual user controls, and enabling multi-grained text editing, it can serve as a new touchstone for LMMs.
+The standard architecture of an LMM comprises a vision backbone for encoding images or videos, a projector (Li et al., 2023b; Cha et al., 2023; Yao et al., 2024) to translate visual embeddings into textual semantic space, and an LLM to process all multimodal contexts.
+As the VCE task involves capturing video semantics, understanding textual user controls, and enabling text editing, it can serve as a new touchstone for LMMs.
 3. Video Caption Editing Task
@@
 We observe that natural language and writing-revision traces are two natural interactive modes.
-The former can be received from keyboards or speech converters, while the latter conveniently expresses user intentions with the prevalence of tablets and wireless stylus pens111https://www.apple.com/apple-pencil/.
+The former can be received from keyboards or speech converters, while the latter conveniently expresses user intentions with the prevalence of tablets and wireless stylus pens.
 A command representation compatible with the above two signals is important and meaningful.
@@
 first delete then add).
-Meanwhile, position and attribute in the triplet are optional, therefore, as shown in Table 1, seven specific commands222Note that we omit the command “del, pos, attr, delete attributes at specified positions”, as it can be covered by “del, pos, -, delete description at specified positions”.
+Meanwhile, position and attribute in the triplet are optional, therefore, as shown in Table 1, seven specific commands111Note that we omit the command “del, pos, attr, delete attributes at specified positions”, as it can be covered by “del, pos, -, delete description at specified positions”.
 via different combinations of operation, position, and attribute elements in the triplet can cover multi-grained realistic demands from coarse-grained (global) controls to fine-grained (local) controls.
@@
 We sample an annotated caption of a video as the reference caption, and the next goal is to construct the command and the related edited caption.
-In general, we aim to construct related (command, edited caption) samples including: 1) coarse-grained length-control commands referring to the global add or delete edits that result in length changes, and 2) finer-grained attribute-related commands to achieve add or delete attributes on the reference caption.
+In general, we aim to construct related (command, edited caption) samples including: 1) coarse-grained length-control commands referring to the global add or delete edits that result in length changes, and 2) finer-grained attribute-related commands to achieve add or delete attributes.
 Coarse-grained length-control commands.
-For add operation, we directly select a longer caption with significant length differences from the rest annotations as the edited caption.
+For add operation, we directly select a longer caption with remarkable length differences as the edited caption.
 The delete operation does the exact opposite.
@@
 We construct attribute samples in a “degradation” manner, that is firstly detecting attribute words in the reference caption and then removing them to obtain the edited caption .
-We utilize Spacy Syntactic Dependency333https://spacy.io/ and Semantic Role Labeling (Shi and Lin, 2019) to achieve noun, verb or modifier attributes detection and removal in while maintaining the fluency of to get .
+We utilize Spacy Syntactic Dependency222https://spacy.io/ and Semantic Role Labeling (Shi and Lin, 2019) to achieve noun, verb or modifier attributes detection and removal in while maintaining the fluency of to get .
 After degradation, we can obtain quadruples for the del, -, attr command.
@@
 We also use the Perplexity (PPL) (Jelinek et al., 1977) metric that reflects the grammatical correctness and semantic meaningfulness.
+| Model | Controllability | Fluency | Vision Align | ||||||
+| Len-Acc | Attr-Acc | Pos-Acc | SARI | BLEU4 | ROUGE-L | PPL | EMScore | ||
+| 1 | Multimodal BART | - | - | - | 49.7 | 48.0 | 62.0 | 73.9 | 28.7 |
+| 2 | 97 | - | - | 52.1 | 49.6 | 63.2 | 70.6 | 28.7 | |
+| 3 | 97 | 93 | - | 53.8 | 52.3 | 65.7 | 72.7 | 28.7 | |
+| 4 | OPA | 98 | 93 | 99 | 53.9 | 52.2 | 65.6 | 72.8 | 28.7 |
+| 5 | 98 | 82 | 97 | 52.6 | 50.5 | 64.4 | 77.2 | 28.7 | |
+| 6 | 3 Single-grained Models | 96 | 69 | 99 | 53.3 | 51.5 | 64.6 | 72.8 | 28.6 |
+Command Controllability Fluency Vision Align Len-Acc Attr-Acc Pos-Acc SARI BLEU4 ROUGE-L PPL EMScore 1 add, - , - 57 - - 26.2 62.1 73.9 23.7 44.7 2 add, pos, - 75 - 59 27.0 83.6 90.3 25.1 44.9 3 add, - , attr 80 74 - 31.7 84.7 89.9 26.6 45.2 4 add, pos, attr 92 70 85 32.9 88.1 93.3 25.5 44.8 5 del, - , - 100 - - 33.5 66.8 73.8 30.5 44.6 6 del, pos, - 99 - - 30.7 83.9 90.6 28.8 44.6 7 del, - , attr 100 93 - 33.6 75.2 83.6 28.9 44.7 8 Overall 84 79 72 30.5 74.4 81.8 26.6 44.8
 Controllability.
@@
 6. Experiments
-| Model | Controllability | Fluency | Vision Align | ||||||
-| Len-Acc | Attr-Acc | Pos-Acc | SARI | BLEU4 | ROUGE-L | PPL | EMScore | ||
-| 1 | Multimodal BART | - | - | - | 49.7 | 48.0 | 62.0 | 73.9 | 28.7 |
-| 2 | 97 | - | - | 52.1 | 49.6 | 63.2 | 70.6 | 28.7 | |
-| 3 | 97 | 93 | - | 53.8 | 52.3 | 65.7 | 72.7 | 28.7 | |
-| 4 | OPA | 98 | 93 | 99 | 53.9 | 52.2 | 65.6 | 72.8 | 28.7 |
-| 5 | 98 | 82 | 97 | 52.6 | 50.5 | 64.4 | 77.2 | 28.7 | |
-| 6 | 3 Single-grained Models | 96 | 69 | 99 | 53.3 | 51.5 | 64.6 | 72.8 | 28.6 |
-Command Controllability Fluency Vision Align Len-Acc Attr-Acc Pos-Acc SARI BLEU4 ROUGE-L PPL EMScore 1 add, - , - 57 - - 26.2 62.1 73.9 23.7 44.7 2 add, pos, - 75 - 59 27.0 83.6 90.3 25.1 44.9 3 add, - , attr 80 74 - 31.7 84.7 89.9 26.6 45.2 4 add, pos, attr 92 70 85 32.9 88.1 93.3 25.5 44.8 5 del, - , - 100 - - 33.5 66.8 73.8 30.5 44.6 6 del, pos, - 99 - - 30.7 83.9 90.6 28.8 44.6 7 del, - , attr 100 93 - 33.6 75.2 83.6 28.9 44.7 8 Overall 84 79 72 30.5 74.4 81.8 26.6 44.8
 6.1. Implementation Details
@@
 For training, we use AdamW (Loshchilov and Hutter, 2017) with a learning rate of 1e-5 and optimize for 20 epochs with a batch size of 20.
-During inference, we set beam size as 5.
+During inference, we set the beam size of generation as 5.
 The ImgLLM pipeline utilizes the identical frame number N as the OPA model.
@@
 In conclusion, the key of the VCE task lies in the combination of fine-grained video and user control understanding and precise text editing capabilities.
+Command Video EMScore SARI BLEU4 Overall ✗ 28.1 53.5 51.7 ✔ 28.7 (+0.6) 53.9 (+0.4) 52.2 (+0.5) del, - , - ✗ 27.0 39.4 12.1 ✔ 28.5 (+1.5) 40.4 (+1.0) 13.5 (+1.4) add, - , - ✗ 28.3 41.9 10.6 ✔ 29.0 (+0.7) 42.0 (+0.1) 11.1 (+0.5)
 6.3. Further Task Analysis
-We conduct further ablation studies on the small-scale OPA model to delve into a detailed analysis of the characteristics of the VCE task.
+We conduct further ablation studies on the small-scale OPA model to delve into a detailed analysis of the VCE task.
 Increase control signals.
@@
 It demonstrates that our unified input design can alleviate the confusion and heterogeneity of multi-grained commands.
+Dataset Model Control. Fluency Vision Align EMMAD-EDIT Trans. 2.94 3.02 3.27 OPA 3.98 3.85 3.76 GT 4.67 4.37 4.22 VATEX-EDIT Trans. 4.18 4.23 3.76 OPA 4.36 4.43 3.93 GT 4.48 4.34 4.41
 Difficulty level of various commands.
@@
 It reveals that when provided with more detailed control signals, the model can generate desired captions more easily.
-Command Video EMScore SARI BLEU4 Overall ✗ 28.1 53.5 51.7 ✔ 28.7 (+0.6) 53.9 (+0.4) 52.2 (+0.5) del, - , - ✗ 27.0 39.4 12.1 ✔ 28.5 (+1.5) 40.4 (+1.0) 13.5 (+1.4) add, - , - ✗ 28.3 41.9 10.6 ✔ 29.0 (+0.7) 42.0 (+0.1) 11.1 (+0.5)
 Effects of vision modality.
@@
 The edited description can serve as the reference caption in the next round to make further editing to satisfy dynamic user demands.
-| Dataset | Model | Control. | Fluency | Vision Align |
-| EMMAD-EDIT | Trans. | 2.94 | 3.02 | 3.27 |
-| OPA | 3.98 | 3.85 | 3.76 | |
-| GT | 4.67 | 4.37 | 4.22 | |
-| VATEX-EDIT | Trans. | 4.18 | 4.23 | 3.76 |
-| OPA | 4.36 | 4.43 | 3.93 | |
-| GT | 4.48 | 4.34 | 4.41 |
 Human Evaluation.
@@
 Further insights into significant future directions are discussed in Appendix F.
+Acknowledgements.
+This work was partially supported by the National Natural Science Foundation of China (No.
+62072462), the Fundamental Research Funds for the Central Universities, and the National Natural Science Foundation of China (No.
+62176002).References
+- (1)
+- Aafaq et al.
+(2019) Nayyer Aafaq, Naveed Akhtar, Wei Liu, Syed Zulqarnain Gilani, and Ajmal Mian.
+2019.
+Spatio-temporal dynamics and semantic attribute enriched visual encoding for video captioning.
+In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition.
+12487–12496.
+- Bai et al.
+(2023) Jinze Bai, Shuai Bai, Shusheng Yang, Shijie Wang, Sinan Tan, Peng Wang, Junyang Lin, Chang Zhou, and Jingren Zhou.
+2023.
+Qwen-vl: A frontier large vision-language model with versatile abilities.
+arXiv preprint arXiv:2308.12966 (2023).
+- Cha et al. (2023) Junbum Cha, Wooyoung Kang, Jonghwan Mun, and Byungseok Roh. 2023. Honeybee: Locality-enhanced Projector for Multimodal LLM. arXiv preprint arXiv:2312.06742 (2023).
+- Chen et al.
+(2019) Shizhe Chen, Qin Jin, Jia Chen, and Alexander G.
+Hauptmann.
+2019.
+Generating Video Descriptions With Latent Topic Guidance.
+IEEE Transactions on Multimedia 21, 9 (2019), 2407–2418.
+https://doi.org/10.1109/TMM.2019.2896515
+- Chowdhery et al.
+(2022) Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, Parker Schuh, Kensen Shi, Sasha Tsvyashchenko, Joshua Maynez, Abhishek Rao, Parker Barnes, Yi Tay, Noam Shazeer, Vinodkumar Prabhakaran, Emily Reif, Nan Du, Ben Hutchinson, Reiner Pope, James Bradbury, Jacob Austin, Michael Isard, Guy Gur-Ari, Pengcheng Yin, Toju Duke, Anselm Levskaya, Sanjay Ghemawat, Sunipa Dev, Henryk Michalewski, Xavier Garcia, Vedant Misra, Kevin Robinson, Liam Fedus, Denny Zhou, Daphne Ippolito, David Luan, Hyeontaek Lim, Barret Zoph, Alexander Spiridonov, Ryan Sepassi, David Dohan, Shivani Agrawal, Mark Omernick, Andrew M.
+Dai, Thanumalayan Sankaranarayana Pillai, Marie Pellat, Aitor Lewkowycz, Erica Moreira, Rewon Child, Oleksandr Polozov, Katherine Lee, Zongwei Zhou, Xuezhi Wang, Brennan Saeta, Mark Diaz, Orhan Firat, Michele Catasta, Jason Wei, Kathy Meier-Hellstern, Douglas Eck, Jeff Dean, Slav Petrov, and Noah Fiedel.
+2022.
+PaLM: Scaling Language Modeling with Pathways.
+CoRR abs/2204.02311 (2022).
+- Dai et al.
+(2023) Wenliang Dai, Junnan Li, Dongxu Li, Anthony Meng Huat Tiong, Junqi Zhao, Weisheng Wang, Boyang Albert Li, Pascale Fung, and Steven C.
+H.
+Hoi.
+2023.
+InstructBLIP: Towards General-purpose Vision-Language Models with Instruction Tuning.
+ArXiv abs/2305.06500 (2023).
+https://api.semanticscholar.org/CorpusID:258615266
+- Deshpande et al. (2019) Aditya Deshpande, Jyoti Aneja, Liwei Wang, Alexander Schwing, and David A Forsyth. 2019. Diverse and controllable image captioning with part-of-speech guidance. CVPR (2019).
+- Dong et al.
+(2022) Qingxiu Dong, Lei Li, Damai Dai, Ce Zheng, Zhiyong Wu, Baobao Chang, Xu Sun, Jingjing Xu, and Zhifang Sui.
+2022.
+A survey for in-context learning.
+arXiv preprint arXiv:2301.00234 (2022).
+- Du et al.
+(2022) Wanyu Du, Vipul Raheja, Dhruv Kumar, Zae Myung Kim, Melissa Lopez, and Dongyeop Kang.
+2022.
+Understanding Iterative Revision from Human-Written Text.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers).
+Association for Computational Linguistics, Dublin, Ireland, 3573–3590.
+https://aclanthology.org/2022.acl-long.250
+- Du (2019) Zeyao Du. 2019. GPT2-Chinese: Tools for training GPT2 model in Chinese language. https://github.com/Morizeyao/GPT2-Chinese.
+- Jelinek et al.
+(1977) Fred Jelinek, Robert L Mercer, Lalit R Bahl, and James K Baker.
+1977.
+Perplexity—a measure of the difficulty of speech recognition tasks.
+The Journal of the Acoustical Society of America 62, S1 (1977), S63–S63.
+- Jin et al.
+(2023) Peng Jin, Ryuichi Takanobu, Caiwan Zhang, Xiaochun Cao, and Li Yuan.
+2023.
+Chat-UniVi: Unified Visual Representation Empowers Large Language Models with Image and Video Understanding.
+arXiv preprint arXiv:2311.08046 (2023).
+- Lewis et al.
+(2020) Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed, Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer.
+2020.
+BART: Denoising Sequence-to-Sequence Pre-training for Natural Language Generation, Translation, and Comprehension.
+In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics.
+Association for Computational Linguistics, Online, 7871–7880.
+https://doi.org/10.18653/v1/2020.acl-main.703
+- Li et al.
+(2023b) Junnan Li, Dongxu Li, Silvio Savarese, and Steven Hoi.
+2023b.
+BLIP-2: Bootstrapping Language-Image Pre-training with Frozen Image Encoders and Large Language Models.
+arXiv:2301.12597 [cs.CV]
+- Li et al.
+(2022a) Junnan Li, Dongxu Li, Caiming Xiong, and Steven Hoi.
+2022a.
+Blip: Bootstrapping language-image pre-training for unified vision-language understanding and generation.
+In International Conference on Machine Learning.
+PMLR, 12888–12900.
+- Li et al.
+(2023a) KunChang Li, Yinan He, Yi Wang, Yizhuo Li, Wenhai Wang, Ping Luo, Yali Wang, Limin Wang, and Yu Qiao.
+2023a.
+Videochat: Chat-centric video understanding.
+arXiv preprint arXiv:2305.06355 (2023).
+- Li et al.
+(2022b) Kunchang Li, Yali Wang, Yinan He, Yizhuo Li, Yi Wang, Limin Wang, and Yu Qiao.
+2022b.
+UniFormerV2: Spatiotemporal Learning by Arming Image ViTs with Video UniFormer.
+arXiv preprint arXiv:2211.09552 (2022).
+- Li et al.
+(2016) Xiangyang Li, Xinhang Song, Luis Herranz, Yaohui Zhu, and Shuqiang Jiang.
+2016.
+Image captioning with both object and scene information.
+In Proceedings of the 24th ACM international conference on Multimedia.
+1107–1110.
+- Lin et al.
+(2023) Bin Lin, Bin Zhu, Yang Ye, Munan Ning, Peng Jin, and Li Yuan.
+2023.
+Video-LLaVA: Learning United Visual Representation by Alignment Before Projection.
+arXiv preprint arXiv:2311.10122 (2023).
+- Lin (2004) Chin-Yew Lin. 2004. Rouge: A package for automatic evaluation of summaries. In Text summarization branches out. 74–81.
+- Liu et al. (2016) Chang Liu, Changhu Wang, Fuchun Sun, and Yong Rui. 2016. Image2Text: a multimodal image captioner. In Proceedings of the 24th ACM international conference on Multimedia. 746–748.
+- Liu et al.
+(2018) Daqing Liu, Zheng-Jun Zha, Hanwang Zhang, Yongdong Zhang, and Feng Wu.
+2018.
+Context-aware visual policy network for sequence-level image captioning.
+In Proceedings of the 26th ACM international conference on Multimedia.
+1416–1424.
+- Liu et al.
+(2021) Fenglin Liu, Xuancheng Ren, Xian Wu, Bang Yang, Shen Ge, and Xu Sun.
+2021.
+O2NA: An Object-Oriented Non-Autoregressive Approach for Controllable Video Captioning.
+In Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021.
+Association for Computational Linguistics, Online, 281–292.
+https://doi.org/10.18653/v1/2021.findings-acl.24
+- Liu et al. (2023) Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. 2023. Visual instruction tuning. arXiv preprint arXiv:2304.08485 (2023).
+- Loshchilov and Hutter (2017) Ilya Loshchilov and Frank Hutter. 2017. Decoupled Weight Decay Regularization. In International Conference on Learning Representations.
+- Luo et al.
+(2020) Huaishao Luo, Lei Ji, Botian Shi, Haoyang Huang, Nan Duan, Tianrui Li, Jason Li, Taroon Bharti, and Ming Zhou.
+2020.
+UniVL: A Unified Video and Language Pre-Training Model for Multimodal Understanding and Generation.
+arXiv preprint arXiv:2002.06353 (2020).
+- Luo et al.
+(2023) Ruipu Luo, Ziwang Zhao, Min Yang, Junwei Dong, Minghui Qiu, Pengcheng Lu, Tao Wang, and Zhongyu Wei.
+2023.
+Valley: Video Assistant with Large Language model Enhanced abilitY.
+arXiv preprint arXiv:2306.07207 (2023).
+- Maaz et al.
+(2023) Muhammad Maaz, Hanoona Rasheed, Salman Khan, and Fahad Shahbaz Khan.
+2023.
+Video-ChatGPT: Towards Detailed Video Understanding via Large Vision and Language Models.
+ArXiv abs/2306.05424 (2023).
+https://api.semanticscholar.org/CorpusID:259108333
+- Müller (2007) Meinard Müller. 2007. Dynamic time warping. Information retrieval for music and motion (2007), 69–84.
+- OpenAI (2022) OpenAI. 2022. OpenAI: Introducing ChatGPT. https://openai.com/blog/chatgpt
+- OpenAI (2023a) OpenAI. 2023a. GPT-4 Technical Report. arXiv:2303.08774 [cs.CL]
+- OpenAI (2023b) OpenAI. 2023b. GPT-4V(ision) System Card. https://cdn.openai.com/papers/GPTV_System_Card.pdf.
+- Ouyang et al.
+(2022) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L.
+Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simens, Amanda Askell, Peter Welinder, Paul F.
+Christiano, Jan Leike, and Ryan Lowe.
+2022.
+Training language models to follow instructions with human feedback.
+CoRR abs/2203.02155 (2022).
+- Pan et al.
+(2020) Boxiao Pan, Haoye Cai, De-An Huang, Kuan-Hui Lee, Adrien Gaidon, Ehsan Adeli, and Juan Carlos Niebles.
+2020.
+Spatio-temporal graph for video captioning with knowledge distillation.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+10870–10879.
+- Papineni et al.
+(2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu.
+2002.
+Bleu: a method for automatic evaluation of machine translation.
+In Proceedings of the 40th annual meeting of the Association for Computational Linguistics.
+311–318.
+- Radford et al.
+(2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, Gretchen Krueger, and Ilya Sutskever.
+2021.
+Learning Transferable Visual Models From Natural Language Supervision.
+In ICML.
+- Ramanishka et al.
+(2016) Vasili Ramanishka, Abir Das, Dong Huk Park, Subhashini Venugopalan, Lisa Anne Hendricks, Marcus Rohrbach, and Kate Saenko.
+2016.
+Multimodal video description.
+In Proceedings of the 24th ACM international conference on Multimedia.
+1092–1096.
+- Ren et al. (2023) Shuhuai Ren, Linli Yao, Shicheng Li, Xu Sun, and Lu Hou. 2023. TimeChat: A Time-sensitive Multimodal Large Language Model for Long Video Understanding. ArXiv abs/2312.02051 (2023).
+- Sammani and Elsayed (2019) Fawaz Sammani and Mahmoud Elsayed. 2019. Look and Modify: Modification Networks for Image Captioning.
+- Sammani and Melas-Kyriazi (2020) Fawaz Sammani and Luke Melas-Kyriazi.
+2020.
+Show, Edit and Tell: A Framework for Editing Image Captions.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR).
+- Shetty and Laaksonen (2016) Rakshith Shetty and Jorma Laaksonen.
+2016.
+Frame-and segment-level features and candidate pool evaluation for video caption generation.
+In Proceedings of the 24th ACM international conference on Multimedia.
+1073–1076.
+- Shi and Lin (2019) Peng Shi and Jimmy Lin. 2019. Simple bert models for relation extraction and semantic role labeling. arXiv preprint arXiv:1904.05255 (2019).
+- Shi et al.
+(2019) Xiangxi Shi, Jianfei Cai, Shafiq Joty, and Jiuxiang Gu.
+2019.
+Watch it twice: Video captioning with a refocused video encoder.
+In Proceedings of the 27th ACM international conference on multimedia.
+818–826.
+- Shi et al.
+(2022) Yaya Shi, Xu Yang, Haiyang Xu, Chunfeng Yuan, Bing Li, Weiming Hu, and Zheng-Jun Zha.
+2022.
+EMScore: Evaluating Video Captioning via Coarse-Grained and Fine-Grained Embedding Matching.
+In IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2022, New Orleans, LA, USA, June 18-24, 2022.
+- Shukang et al.
+(2023) Yin Shukang, Fu Chaoyou, Zhao Sirui, Xu Tong, Wang Hao, Sui Dianbo, Shen Yunhang, Li Ke, Sun Xing, and Chen Enhong.
+2023.
+Woodpecker: Hallucination Correction for Multimodal Large Language Models.
+arXiv preprint arXiv:2310.16045 (2023).
+- Stefanini et al.
+(2022) Matteo Stefanini, Marcella Cornia, Lorenzo Baraldi, Silvia Cascianelli, Giuseppe Fiameni, and Rita Cucchiara.
+2022.
+From show to tell: a survey on deep learning-based image captioning.
+IEEE transactions on pattern analysis and machine intelligence 45, 1 (2022), 539–559.
+- Sun et al. (2023) Quan Sun, Yuxin Fang, Ledell Wu, Xinlong Wang, and Yue Cao. 2023. Eva-clip: Improved training techniques for clip at scale. arXiv preprint arXiv:2303.15389 (2023).
+- Tang et al.
+(2021) Mingkang Tang, Zhanyu Wang, Zhenhua Liu, Fengyun Rao, Dian Li, and Xiu Li.
+2021.
+Clip4caption: Clip for video caption.
+In Proceedings of the 29th ACM International Conference on Multimedia.
+4858–4862.
+- Touvron et al.
+(2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, Aurélien Rodriguez, Armand Joulin, Edouard Grave, and Guillaume Lample.
+2023a.
+LLaMA: Open and Efficient Foundation Language Models.
+CoRR (2023).
+- Touvron et al.
+(2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al.
+2023b.
+Llama 2: Open foundation and fine-tuned chat models.
+arXiv preprint arXiv:2307.09288 (2023).
+- Venugopalan et al.
+(2015) Subhashini Venugopalan, Marcus Rohrbach, Jeffrey Donahue, Raymond Mooney, Trevor Darrell, and Kate Saenko.
+2015.
+Sequence to Sequence – Video to Text.
+In 2015 IEEE International Conference on Computer Vision (ICCV).
+4534–4542.
+https://doi.org/10.1109/ICCV.2015.515
+- Vinyals et al.
+(2015) Oriol Vinyals, Alexander Toshev, Samy Bengio, and Dumitru Erhan.
+2015.
+Show and tell: A neural image caption generator.
+In Proceedings of the IEEE conference on computer vision and pattern recognition.
+3156–3164.
+- Wang et al.
+(2019a) Bairui Wang, Lin Ma, Wei Zhang, Wenhao Jiang, Jingwen Wang, and Wei Liu.
+2019a.
+Controllable Video Captioning with POS Sequence Guidance Based on Gated Fusion Network.
+arXiv preprint arXiv:1908.10072 (2019).
+- Wang et al.
+(2019b) Xin Wang, Jiawei Wu, Junkun Chen, Lei Li, Yuan-Fang Wang, and William Yang Wang.
+2019b.
+Vatex: A large-scale, high-quality multilingual dataset for video-and-language research.
+In Proceedings of the IEEE/CVF International Conference on Computer Vision.
+4581–4591.
+- Wang et al.
+(2022) Zhen Wang, Long Chen, Wenbo Ma, Guangxing Han, Yulei Niu, Jian Shao, and Jun Xiao.
+2022.
+Explicit Image Caption Editing.
+In Computer Vision–ECCV 2022: 17th European Conference, Tel Aviv, Israel, October 23–27, 2022, Proceedings, Part XXXVI.
+113–129.
+- Wolf et al.
+(2020) Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, Rémi Louf, Morgan Funtowicz, et al.
+2020.
+Transformers: State-of-the-art natural language processing.
+In Proceedings of the 2020 conference on empirical methods in natural language processing: system demonstrations.
+38–45.
+- Xu et al.
+(2015) Kelvin Xu, Jimmy Ba, Ryan Kiros, Kyunghyun Cho, Aaron Courville, Ruslan Salakhudinov, Rich Zemel, and Yoshua Bengio.
+2015.
+Show, attend and tell: Neural image caption generation with visual attention.
+In International conference on machine learning.
+PMLR, 2048–2057.
+- Xu et al.
+(2016) Wei Xu, Courtney Napoles, Ellie Pavlick, Quanze Chen, and Chris Callison-Burch.
+2016.
+Optimizing Statistical Machine Translation for Text Simplification.
+Transactions of the Association for Computational Linguistics 4 (2016), 401–415.
+https://doi.org/10.1162/tacl_a_00107
+- Yang et al.
+(2022) An Yang, Junshu Pan, Junyang Lin, Rui Men, Yichang Zhang, Jingren Zhou, and Chang Zhou.
+2022.
+Chinese CLIP: Contrastive Vision-Language Pretraining in Chinese.
+arXiv preprint arXiv:2211.01335 (2022).
+- Yang et al.
+(2023a) Dingyi Yang, Hongyu Chen, Xinglin Hou, Tiezheng Ge, Yuning Jiang, and Qin Jin.
+2023a.
+Visual captioning at will: Describing images and videos guided by a few stylized sentences.
+In Proceedings of the 31st ACM International Conference on Multimedia.
+5705–5715.
+- Yang et al.
+(2023b) Zhengyuan Yang, Linjie Li, Kevin Lin, Jianfeng Wang, Chung-Ching Lin, Zicheng Liu, and Lijuan Wang.
+2023b.
+The Dawn of LMMs: Preliminary Explorations with GPT-4V(ision).
+arXiv:2309.17421 [cs.CV]
+- Yao et al. (2023) Linli Yao, Weijing Chen, and Qin Jin. 2023. CapEnrich: Enriching Caption Semantics for Web Images via Cross-modal Pre-trained Knowledge. In TheWebConf.
+- Yao et al.
+(2024) Linli Yao, Lei Li, Shuhuai Ren, Lean Wang, Yuanxin Liu, Xu Sun, and Lu Hou.
+2024.
+DeCo: Decoupling Token Compression from Semantic Abstraction in Multimodal Large Language Models.
+arXiv preprint arXiv:2405.20985 (2024).
+- Yao et al.
+(2022) Linli Yao, Weiying Wang, and Qin Jin.
+2022.
+Image difference captioning with pre-training and contrastive learning.
+In Proceedings of the AAAI Conference on Artificial Intelligence, Vol.
+36.
+3108–3116.
+- Ye et al.
+(2023a) Qinghao Ye, Haiyang Xu, Guohai Xu, Jiabo Ye, Ming Yan, Yiyang Zhou, Junyang Wang, Anwen Hu, Pengcheng Shi, Yaya Shi, et al.
+2023a.
+mplug-owl: Modularization empowers large language models with multimodality.
+arXiv preprint arXiv:2304.14178 (2023).
+- Ye et al.
+(2023b) Qinghao Ye, Haiyang Xu, Jiabo Ye, Ming Yan, Anwen Hu, Haowei Liu, Qi Qian, Ji Zhang, Fei Huang, and Jingren Zhou.
+2023b.
+mPLUG-Owl2: Revolutionizing Multi-modal Large Language Model with Modality Collaboration.
+arXiv:2311.04257 [cs.CL]
+- Yu et al.
+(2017) Youngjae Yu, Hyungjin Ko, Jongwook Choi, and Gunhee Kim.
+2017.
+End-to-End Concept Word Detection for Video Captioning, Retrieval, and Question Answering.
+In 2017 IEEE Conference on Computer Vision and Pattern Recognition (CVPR).
+3261–3269.
+https://doi.org/10.1109/CVPR.2017.347
+- Yuan et al.
+(2023) Mengqi Yuan, Bing-Kun Bao, Zhiyi Tan, and Changsheng Xu.
+2023.
+Adaptive Text Denoising Network for Image Caption Editing.
+ACM Trans.
+Multimedia Comput.
+Commun.
+Appl.
+19, 1s, Article 41 (feb 2023), 18 pages.
+https://doi.org/10.1145/3532627
+- Yuan et al. (2020) Yitian Yuan, Lin Ma, Jingwen Wang, and Wenwu Zhu. 2020. Controllable Video Captioning with an Exemplar Sentence. In the 28th ACM International Conference on Multimedia (MM ’20).
+- Zeng et al.
+(2022) Aohan Zeng, Xiao Liu, Zhengxiao Du, Zihan Wang, Hanyu Lai, Ming Ding, Zhuoyi Yang, Yifan Xu, Wendi Zheng, Xiao Xia, Weng Lam Tam, Zixuan Ma, Yufei Xue, Jidong Zhai, Wenguang Chen, Peng Zhang, Yuxiao Dong, and Jie Tang.
+2022.
+GLM-130B: An Open Bilingual Pre-trained Model.
+abs/2210.02414 (2022).
+- Zhang et al. (2023) Hang Zhang, Xin Li, and Lidong Bing. 2023. Video-llama: An instruction-tuned audio-visual language model for video understanding. arXiv preprint arXiv:2306.02858 (2023).
+- Zhang et al.
+(2022) Zhipeng Zhang, Xinglin Hou, Kai Niu, Zhongzhen Huang, Tiezheng Ge, Yuning Jiang, Qi Wu, and Peng Wang.
+2022.
+Attract me to Buy: Advertisement Copywriting Generation with Multimodal Multi-structured Information.
+arXiv preprint arXiv:2205.03534 (2022).
+In the supplementary material, we first introduce details of three approaches (Section A), then present more construction details about the VATEX-EDIT dataset (Section B) and EMMAD-EDIT dataset (Section C).
+Moreover, we provide the pseudo-code of the Position Accuracy metric (Section D) and the conversion instructions from interface signals to triplet format (Section E).
+Finally, based on a good start on the task, dataset and method foundation, we further discuss a variety of interesting aspects worth exploring in the future (Section F) and the related social impact (Section G).
 Appendix A Model Details
 A.1. Architecture of OPA Model
-The overall architecture of our specialist model, i.e. OPA, is depicted in Figure 8.
+The overall architecture of the specialist model OPA is depicted in Figure 8, which is built on the BART model.
 A.2. Prompt of ImgLLM Pipeline
@@
 Figure 9 illustrates the converted data sample.
+A.4. Results of GPT4-o
+As Table 7 shows, we report the results of GPT-4o333https://openai.com/index/hello-gpt-4o/ in the challenging EMMAD-EDIT dataset as a reference upper bound of the task.
+Given that the video interface for GPT-4o is not yet available, we extract 8 frames from each video and assess the model using these multiple images as inputs.
+Control. Fluency Text-Vision Align OPA 66.4 76.5 44.8 VidLLM (VideoChat) 62.6 81.7 45.3 ImgLLM (w/ ChatGPT) 66.9 79.8 45.3 GPT-4o 71.1 81.9 45.3
 Appendix B VATEX-EDIT Dataset Details
@@
 We construct quadruples (video, command, reference caption, edited caption) according to two types of commands including coarse-grained length-control commands and fine-grained attribute-related commands.
-The overall procedure is depicted in Figure 11.
 We complement more details about constructing attribute-related commands.
@@
 In detail, we adopt four steps as follows:
-First, we use the Spacy444https://spacy.io/ syntactic dependency parser to build a textual dependency tree that contains the Part-of-Speech information and relationships between tokens.
+First, we use the Spacy syntactic dependency parser to build a textual dependency tree that contains the Part-of-Speech information and relationships between tokens.
 We select reasonable branches in the parsed tree as attributes and further prune the branch to ensure the fluency of the rest caption.
@@
 In detail, we replace the edited caption by retrieving desired sentences satisfying both length and semantics (attributes) constraints from original annotated descriptions for the same video.
-| Constructed | Annotated | |
-| ChatGPT | 87% | 91% |
-| Human | 4.40 | 4.37 |
 B.2. VATEX-EDIT Test set Quality
@@
 To assess the quality of the auto-constructed sentences, we evaluate the test set of VATEX-EDIT through human and ChatGPT evaluations.
-For the ChatGPT555https://openai.com/blog/chatgpt evaluation, we design suitable prompts (depicted in Figure 12) to guide ChatGPT to judge whether a sentence is fluent or not.
+For the ChatGPT evaluation, we design suitable prompts (depicted in Figure 12) to guide ChatGPT to judge whether a sentence is fluent or not.
 We randomly sample 100 ground-truth sentences on the test set and calculate the fluency rate obtained by ChatGPT.
@@
 As Table 8 shows, the automatically constructed ground-truth sentences of the “del, pos, -” command are as fluent as the human-annotated ground-truth sentences of other commands (87% vs 91% and 4.40 vs 4.37), which indicates the high quality of the VATEX-EDIT test set.
+Appendix C EMMAD-EDIT Dataset Details
+We manually collect the high-quality dataset EMMAD-EDIT in the E-commerce domain based on the Chinese E-commerce video captioning dataset E-MMAD.
+The data sample in E-MMAD dataset consists of a product video, an advertising video description, and additional information including video titles and structure attributes.
+It collects 120,984 videos with average duration of 30.4 seconds and the annotated Chinese description length is 67 words on average.
+The characteristics of long videos and long captions make it suitable for building a challenging VCE dataset.
+During annotation, we select data samples from E-MMAD dataset with relatively long descriptions.
+In addition to videos and descriptions, we also provide product structure information and video titles for reference.
+| Constructed | Annotated | |
+| ChatGPT | 87% | 91% |
+| Human | 4.40 | 4.37 |
+C.1. EMMAD-EDIT dataset statistics
+Table 9 shows breakdown data statistics of the EMMAD-EDIT dataset.
+It has overall 79,652 editing instances for 16,176 product videos.
+Diverse unique attributes and vocabulary also indicate data richness.
+We separate the abstract attribute-related data samples as an extra challenging subset.
+Considering realistic demands, we utilize all these data samples to construct “add, -, attr, add attributes in description” command cases.
+Note that in the Experiments section, we present the EMMAD-EDIT results training without the abstract subset data by default.
 | Command | #Videos | #Editing instances | VTime | Edit Dist | Uni. Attrs | Vocab | ||||||
@@
 | 15,955 | 5,328 | 5,432 | 15,959 | 5,328 | 5,432 | 26.8 | 90.9 | 100.9 | 11.4 | 648 | 44,347 | |
-Appendix C EMMAD-EDIT Dataset Details
-We manually collect the high-quality dataset EMMAD-EDIT in the E-commerce domain based on the Chinese E-commerce video captioning dataset E-MMAD.
-The data sample in E-MMAD dataset consists of a product video, an advertising video description, and additional information including video titles and structure attributes.
-It collects 120,984 videos with average duration of 30.4 seconds and the annotated Chinese description length is 67 words on average.
-The characteristics of long videos and long captions make it suitable for building a challenging VCE dataset.
-During annotation, we select data samples from E-MMAD dataset with relatively long descriptions.
-In addition to videos and descriptions, we also provide product structure information and video titles for reference.
-C.1. EMMAD-EDIT dataset statistics
-Table 8 shows breakdown data statistics of the EMMAD-EDIT dataset.
-It has overall 79,652 editing instances for 16,176 product videos.
-Diverse unique attributes and vocabulary also indicate data richness.
-We separate the abstract attribute-related data samples as an extra challenging subset.
-Considering realistic demands, we utilize all these data samples to construct “add, -, attr, add attributes in description” command cases.
-Note that in the Experiments section, we present the EMMAD-EDIT results training without the abstract subset data by default.
 C.2. The Impact of Data Volume.
@@
 To tackle the above challenge, we propose a novel Dynamic Sequence Aligning (DSA) algorithm to align two variable-length textual sequences based on the absolute positions, inspired by classical Dynamic Time Warping (DTW) (Müller, 2007).
-The POS-Acc metric can align two variable-length text sequences in positional indexes, resulting in related spans {} in aligned to [MASK] tokens in .
+The pseudo-code is presented in Figure 17, which can align two variable-length text sequences in positional indexes, resulting in related spans {} in aligned to [MASK] tokens in .
 We count the percentage of correct samples that insert/remove new content in given position .
@@
 The triplet command in our annotated datasets can be easily converted to natural languages to support more diverse scenarios and applications.
-For example, using natural language to explore the capability of LLMs in the VCE task (A.2).
+For example, using natural language to explore the capability of LLMs in the VCE task (A).
 Specifically, we design a series of templates shown in Table 10 to convert the triplet command to natural language.
 We will also release the two benchmark datasets with user commands both in the triplet and natural language format to benefit the community.
-| Command | Conversion Template |
-| del, - , attr | Delete contents about ’{}’ from the reference description. |
-| add, pos, attr | Add contents about ’{}’ at [POS]. |
-| add, - , attr | Add contents about ’{}’ to expand the reference description. |
-| add, pos, - | Add video-related contents at [POS]. |
-| del, pos, - | Contents have been deleted at [POS], please make the rest sentence fluent and coherent. |
-| del, - , - | Shorten the reference description. |
-| del, - , attr | Expand the reference description. |
 E.3. From Handwriting-revision Traces to Triplets
@@
 The VCE benchmark can serve as a new multi-modal evaluation for assessing both long video understanding and multi-grained text editing abilities.
+| Command | Conversion Template |
+| del, - , attr | Delete contents about ’{}’ from the reference description. |
+| add, pos, attr | Add contents about ’{}’ at [POS]. |
+| add, - , attr | Add contents about ’{}’ to expand the reference description. |
+| add, pos, - | Add video-related contents at [POS]. |
+| del, pos, - | Contents have been deleted at [POS], please make the rest sentence fluent and coherent. |
+| del, - , - | Shorten the reference description. |
+| del, - , attr | Expand the reference description. |
+Appendix G Social Impacts
+The proposed Video Caption Editing (VCE) task addresses significant challenges in the realm of video content management, offering far-reaching social implications.
+Enhanced Accessibility.
+By allowing for multi-grained user control over video captions, our system can significantly improve accessibility for individuals with disabilities.
+Users who rely on captions for understanding video content, such as the hearing-impaired, can benefit from more precise and customizable descriptions tailored to their specific needs.
+E-commerce and Marketing.
+The integration of our VCE system into e-commerce platforms can revolutionize how product videos are presented.
+By enabling dynamic and detailed video descriptions that cater to individual consumer preferences, businesses can enhance user engagement and improve product understanding.
+User Empowerment and Creativity.
+The ability to edit video captions interactively empowers users to tailor content to their specific and creative desires.
+This democratizes content creation, allowing individuals and small creators to produce high-quality, customized video content without requiring extensive technical expertise.